OpenStudy (allieatthedisco):

The coordinates of point T are given. The midpoint of line ST is (5,-8). Find the coordinate of point S if T is (5,-15)

OpenStudy (anonymous):

Let S ( x, y) ( x + 5 ) / 2 = 5 ( y - 15) / 2 = -8 Can you find x, y now?

OpenStudy (allieatthedisco):

would i multiply the 5 in (x+5) by the 2 to find X? and the same process for Y?

OpenStudy (anonymous):

Yes, you're right, that's how you eliminate the fraction :)

OpenStudy (allieatthedisco):

ok, thank you! :)

OpenStudy (anonymous):

What's your result x, y?

OpenStudy (allieatthedisco):

x = 10, y = 30?

OpenStudy (anonymous):

How???

OpenStudy (allieatthedisco):

5*2=10, 15*2=30

OpenStudy (anonymous):

That's why it's necessary to post your result: ( x + 5 ) / 2 = 5 -> x + 5 = 5 * 2 => x = 10 - 5 = 5

OpenStudy (anonymous):

( y - 15) / 2 = -8 -> y - 15 = -8 * 2 .... Can you continue to show how much you understand this time?

OpenStudy (allieatthedisco):

ok, so for this problem would my answer be (0,7) or (5,-1)?

OpenStudy (anonymous):

Where do you get ( 0, 7) :0

OpenStudy (allieatthedisco):

i actually called this local homework hotline place called askrose and the guy tried explaining to to me there and said something like since T = (5,-15) and line ST = (5,-8) you move 0 places for the 5's and move 7 places to get from -15 to -8 thus equaling (0,7)

OpenStudy (anonymous):

I don't know any guy from local hotline, but it's very confusing to explain Math through phone call! All you do is to follow the steps that I've showed you, ask any questions that bother you, so you learn one time for all, because these are fundamental concept which extremely essential for your learning :)

OpenStudy (allieatthedisco):

yes, it was quite hard to follow what he was saying! if i do have any further questions, i will post on here again! thank you :)

OpenStudy (anonymous):

Okie, just continue to practice then you'll surely feel more and more confident :)